I need help in this rational expression
(x^2+2x) divide (x+2)
(x^2-4) (x-2)

Answer by HyperBrain(694)

x^2+2x=x(x+2)
So, (x^2+2x)/(x+2)=x
---
(x^2-4)=(x+2)(x-2)
So, (x^2-4)/(x-2)=x+2